OpenStudy (anonymous):

I know is not the right topic but if you help me i will give you a medal even fan you please help me! Im not being lazy i just dont uderstand! C = (°F - 32) 5/9 ; °F = °C 9/5 + 32; °K = °C + 273; °K = 5/9 (°F - 32) + 273

OpenStudy (anonymous):

sorry this is the riNote: You may use the following formulas to help you complete the data table: °C = (°F - 32) 5/9 ; °F = °C 9/5 + 32; °K = °C + 273; °K = 5/9 (°F - 32) + 273 Planet Name Average Temperature (°F) Average Temperature (°C) Average Temperature (°K) Mercury Venus Earth Mars Jupiter Saturn Uranus Neptune ght thing

OpenStudy (anonymous):

@Jhannybean @Kitt020912 @LazyBoy @ketz someone please help me